# Bogomolov property and Galois representations

### v3

@@ 2. Notations and preliminary results
and “res” denotes the restriction. Let $\overline{H_{n}}$ be the normal closure of $H_{n}$ in $\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$. Recall that $\overline{H_{n}}\subseteq\Omega_{n}$. We want to show that $\mathcal{L}_{n}(\overline{H_{n}})=\widehat{M_{2}}(\mathcal{O}/p\mathcal{O})$. This will imply that $\overline{H_{n}}=\Omega_{n}$ coincides with $\Omega_{n}$ as we want (and that $\mathcal{L}_{n}$ is surjective). Note that, by property 1.3 $(P2)$ $\bar{\rho}_{1}$ is surjective. Since “res” is obviously surjective, $\tilde{\rho}_{n}$ is surjective as well. Thus (5.1) says in particular that $\mathcal{L}_{n}(\overline{H_{n}})$ is invariant by the adjoint action [^5] That is, the action by conjugation. of $\widehat{G}(\mathcal{O}/p\mathcal{O})$. Note that $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ is also invariant by this action. Thus
@@
Indeed, $\mathcal{L}_{n}(H_{n})$ is contained in $\mathrm{M}_{2}({\mathbb{F}}_{p})$ by Proposition 4.2; since $|H_{n}|=p^{2}$ by Lemma 4.7(i), its intersection with $\mathrm{M}_{2}({\mathbb{F}}_{p})^{0}$ is already not trivial.
Then, by Proposition A.2.5, $V=\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ and thus $\mathcal{L}_{n}(\overline{H_{n}})$ contains $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$.
@@
$\rho_{v}|_{G_{\mathbb{Q}_{q}}}\simeq(\epsilon_{2}\oplus\epsilon^{\prime}_{2})\otimes\mu_{\sqrt{-1}}|_{G_{\mathbb{Q}_{q}}}$
(see the proof of Lemma 4.7), its image coincides with $N_{\mathbb{Q}_{q}/\mathbb{Q}_{p}}(1+p^{n-1}\mathbb{Z}_{q})$ mod $1+p^{n}\mathbb{Z}_{p}$, and the norm map from $1+p^{n-1}\mathbb{Z}_{q}\rightarrow 1+p^{n-1}\mathbb{Z}_{p}$ is surjective (see for example [26, Ch. V, §2, Prop. 3]). By condition 1.3 $(P3)$ $p\nmid k-1$ and thus the $(k-1)$-th power of the cyclotomic character is surjective. Therefore also the trace from $\mathcal{L}_{n}(\overline{H_{n}})$ to ${\mathbb{F}}_{p}$ is surjective.
Thus $\mathcal{L}_{n}(\overline{H_{n}})=\widehat{M_{2}}(\mathcal{O}/p\mathcal{O})$ as claimed.
@@
We want to apply Proposition 3.4 with ${\mathbb{L}}={\mathbb{Q}}(p^{\infty})$. We first check Assumption 3.2. By Lemma 4.7 (iii) (with $M=2$) there exists an element $\tau\in\mathrm{Gal}(\mathbb{Q}_{q}(p^{\infty})/\mathbb{Q}_{q})$ whose image by $\rho_{v}$ is the scalar matrix $\left(\begin{smallmatrix}2^{k-1}&0\\ 0&2^{k-1}\end{smallmatrix}\right)$. Since $\rho_{v}|_{G_{\mathbb{Q}_{q}}}$ is totally ramified and $\mathbb{Q}_{q}/\mathbb{Q}_{p}$ is unramified, $\tau$ can be lifted to an element in the inertia subgroup of $G_{\mathbb{Q}_{p}}$. As its image is a scalar, $\tau$ lies in the center of $\mathrm{Gal}({\mathbb{L}}/{\mathbb{Q}})$. Its action on the $p^{\infty}$-th roots of unity is given by the $p$-adic cyclotomic character, i.e. the determinant of $\rho_{v}(\tau)$, which is $4^{k-1}$. Then $\tau(\zeta)=\zeta^{4^{k-1}}$ for every $\zeta\in{\boldsymbol{\mu}}_{p^{\infty}}$.
We now check Assumption 3.3. We consider the filtration given by ${\mathbb{L}}_{n}={\mathbb{Q}}(p^{n})$. Let $n\geq 1$. Fix a place of ${\mathbb{Q}}(p^{\infty})$ over $p$ corresponding to an embedding ¯¯$\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}}_{p}$. Then, by Lemma 4.7 $(i)$ and $(ii)$, and in the notation of Assumption (3.3), $e_{n}=\delta q^{n-1}=\frac{(q-1)q^{n-1}}{(q-1,k-1)}$, $i_{n}=\frac{q^{n-1}-1}{(q-1,k-1)}$ and $H_{n}\simeq\mathrm{Gal}(\mathbb{Q}_{q}(p^{n})/\mathbb{Q}_{q}(p^{n-1}))$. Thus
@@
${\mathbb{F}}(\rho_{p})=\mathbb{Q}(\rho^{(p)}\times\rho_{p})=\mathbb{Q}(\rho).$
@@ A.1. A result on the normal closure
$\widehat{G}({\mathcal{A}})=\{\gamma\in\mathrm{GL}_{2}({\mathcal{A}})\ |\ \det(\gamma)\in(\mathbb{F}_{p}^{\times})^{k-1}\}$
It is the semidirect product $\mathrm{SL}_{2}({\mathcal{A}})\rtimes(\mathbb{F}_{p}^{\times})^{k-1}$, where $x\in(\mathbb{F}^{\times})^{k-1}$ acts on $\mathrm{SL}_{2}({\mathcal{A}})$ by conjugation by $\left(\begin{smallmatrix}x&0\\ 0&1\end{smallmatrix}\right)$. The aim of this subsection is to prove the following proposition.
@@
Let $N$ be the normal closure of $H$ in $\widehat{G}(\mathbb{F}_{p})$. Since $|N|>2s$, the intersection $N\cap\mathrm{SL}_{2}(\mathbb{F}_{p})$ is a non trivial normal subgroup of $\mathrm{SL}_{2}(\mathbb{F}_{p})$ of order strictly greater than $2$. Since $|\mathbb{F}|\geq 4$, the only proper normal subgroup of $SL_{2}(\mathbb{F})$ is {±$\{\pm Id\}$ (see for example [29, Theorem 8.3 and Lemma 8.6, p.539]). Thus $\mathrm{SL}_{2}(\mathbb{F}_{p})\subseteq N$. Since the projection over $(\mathbb{F}_{p}^{\times})^{k-1}$ is surjective, the result follows. ∎
@@ Acknowledgements
We are grateful to Andrea Conti, Samuel Le Fourn and Davide Lombardi for helpful discussions.
We thank the INdAM GNSAGA for support. We acknowledge the IEA (International Emerging Action) project PAriAIPP (Problèmes sur l’Arithmétique et l’Algèbre des Petits Points) of the CNRS for the financial support.

### v4

@@ 2. Notations and preliminary results
and “res” denotes the restriction. Let $\overline{H_{n}}$ be the normal closure of $H_{n}$ in $\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$. Recall that $\overline{H_{n}}\subseteq\Omega_{n}$. We want to show that $\mathcal{L}_{n}(\overline{H_{n}})=\widehat{M_{2}}(\mathcal{O}/p\mathcal{O})$. This will imply that $\overline{H_{n}}=\Omega_{n}$ coincides with $\Omega_{n}$ as we want (and that $\mathcal{L}_{n}$ is surjective). Note that, by Assumption 1.3 $(P2)$ $\bar{\rho}_{1}$ is surjective. Since “res” is obviously surjective, $\tilde{\rho}_{n}$ is surjective as well. Thus (5.1) says in particular that $\mathcal{L}_{n}(\overline{H_{n}})$ is invariant by the adjoint action [^5] That is, the action by conjugation. of $\widehat{G}(\mathcal{O}/p\mathcal{O})$. Note that $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ is also invariant by this action. Thus
@@
Indeed, $\mathcal{L}_{n}(H_{n})$ is contained in $\mathrm{M}_{2}({\mathbb{F}}_{p})$ by Proposition 4.2; since $|H_{n}|=p^{2}$ by Proposition 4.7$(i)$, its intersection with $\mathrm{M}_{2}({\mathbb{F}}_{p})^{0}$ is already not trivial. Then, by Proposition A.2.5, $V=\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ and thus $\mathcal{L}_{n}(\overline{H_{n}})$ contains $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$.
@@
$\rho_{v}|_{G_{\mathbb{Q}_{q}}}\simeq(\varepsilon_{2}\oplus\varepsilon^{\prime}_{2})\otimes\mu_{\sqrt{-1}}|_{G_{\mathbb{Q}_{q}}}$
(see the proof of Proposition 4.7), its image coincides with $N_{\mathbb{Q}_{q}/\mathbb{Q}_{p}}(1+p^{n-1}\mathbb{Z}_{q})$ mod $1+p^{n}\mathbb{Z}_{p}$, and the norm map from $1+p^{n-1}\mathbb{Z}_{q}\rightarrow 1+p^{n-1}\mathbb{Z}_{p}$ is surjective (see for example [27, Ch. V, §2, Prop. 3]). By Assumption 1.3 $(P3)$ $p\nmid k-1$ and thus the $(k-1)$-th power of the cyclotomic character is surjective. Therefore also the trace from $\mathcal{L}_{n}(\overline{H_{n}})$ to ${\mathbb{F}}_{p}$ is surjective.
Thus $\mathcal{L}_{n}(\overline{H_{n}})=\widehat{M_{2}}(\mathcal{O}/p\mathcal{O})$ as claimed. ∎
@@
We want to apply Proposition 3.4 with ${\mathbb{L}}={\mathbb{Q}}(p^{\infty})$. We first check Assumption 3.2. By Proposition 4.7 $(iii)$ (with $M=2$) there exists an element $\tau\in\mathrm{Gal}(\mathbb{Q}_{q}(p^{\infty})/\mathbb{Q}_{q})$ whose image by $\rho_{v}$ is the scalar matrix $\left(\begin{smallmatrix}2^{k-1}&0\\ 0&2^{k-1}\end{smallmatrix}\right)$. Since $\rho_{v}|_{G_{\mathbb{Q}_{q}}}$ is totally ramified and $\mathbb{Q}_{q}/\mathbb{Q}_{p}$ is unramified, $\tau$ can be lifted to an element in the inertia subgroup of $G_{\mathbb{Q}_{p}}$. As its image is a scalar, $\tau$ lies in the center of $\mathrm{Gal}({\mathbb{L}}/{\mathbb{Q}})$. Its action on the $p^{\infty}$-th roots of unity is given by the $p$-adic cyclotomic character $\varepsilon$, and $4^{k-1}=\det(\rho_{v}(\tau))=\varepsilon(\tau)^{k-1}$. Then ${\epsilon(\tau)}=4\omega$, where $\omega$ is a root of unity in $\mathbb{Z}_{p}^{\times}$. Up to replace $\tau$ by $\tau^{p-1}$ we obtain an element that meets the required conditions.
We now check Assumption 3.3. We consider the filtration given by ${\mathbb{L}}_{n}={\mathbb{Q}}(p^{n})$. Let $n\geq 1$. Fix a place of ${\mathbb{Q}}(p^{\infty})$ over $p$ corresponding to an embedding ¯¯$\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}}_{p}$. Then, by Proposition 4.7 $(i)$ and $(ii)$, and in the notation of Assumption (3.3), $e_{n}=\delta q^{n-1}=\frac{(q-1)q^{n-1}}{(q-1,k-1)}$, $i_{n}=\frac{q^{n-1}-1}{(q-1,k-1)}$ and $H_{n}\simeq\mathrm{Gal}(\mathbb{Q}_{q}(p^{n})/\mathbb{Q}_{q}(p^{n-1}))$. Thus
@@
${\mathbb{F}}(\rho_{p})=\mathbb{Q}(\rho^{(p)}\times\rho_{p})=\mathbb{Q}(\rho).\qed$
For a fixed $n\geq 0$, and a field $\mathbb{F}\subseteq\overline{\mathbb{Q}}$, consider the condition
(6.1) $\mathbb{F}$ is Galois, unramified at $p$ and $\mathrm{Frob}_{p}^{n}$ is central in $\mathrm{Gal}(\mathbb{F}/{\mathbb{Q}})$.
It is preserved under composition of fields. Let $\widetilde{\mathbb{F}}_{p,n}$ be the maximal extension satisfying (6.1). The field $\widetilde{\mathbb{F}}_{p,n}$ has (B), by Proposition 3.1, and $\widetilde{\mathbb{F}}_{p,n}\subseteq\widetilde{\mathbb{F}}_{p,m}$ if $n\mid m$. If $f$ fulfils Assumption 1.3, then $\mathbb{Q}(\rho^{(p)})\subseteq\widetilde{\mathbb{F}}_{p,2}$, so that, by Proposition 3.4, Theorem 1.5 has the following
Corollary 6.2.
Let $f=\sum_{n\geq 1}a_{n}q^{n}\in S_{k}(\Gamma_{0}(N))$ be a normalized eigenform. Assume that there exists a rational prime $p$ satisfying Assumption 1.3. Then for every $n\in\mathbb{N}$, the field $\widetilde{\mathbb{F}}_{p,n}(\rho_{f})$ has property (B).
Notice that $\widetilde{\mathbb{F}}_{p,1}$ contains $\mathbb{Q}^{\mathrm{tp}}$, the maximal totally $p$-adic extension of $\mathbb{Q}$. Therefore, when $f$ is a modular form arising from an elliptic curve defined over $\mathbb{Q}$, our Corollary 6.2 intersects to some extent with Theorem 1.6 of [20].
@@ A.1. A result on the normal closure
$\widehat{G}({\mathcal{A}})=\{\gamma\in\mathrm{GL}_{2}({\mathcal{A}})\ |\ \det(\gamma)\in(\mathbb{F}_{p}^{\times})^{k-1}\}.$
It is the semidirect product $\mathrm{SL}_{2}({\mathcal{A}})\rtimes(\mathbb{F}_{p}^{\times})^{k-1}$, where $x\in(\mathbb{F}_{p}^{\times})^{k-1}$ acts on $\mathrm{SL}_{2}({\mathcal{A}})$ by conjugation by $\left(\begin{smallmatrix}x&0\\ 0&1\end{smallmatrix}\right)$. The aim of this subsection is to prove the following proposition.
@@
Let $N$ be the normal closure of $H$ in $\widehat{G}(\mathbb{F}_{p})$. Since $|N|>2s$, the intersection $N\cap\mathrm{SL}_{2}(\mathbb{F}_{p})$ is a non trivial normal subgroup of $\mathrm{SL}_{2}(\mathbb{F}_{p})$ of order strictly greater than $2$. Since $|\mathbb{F}_{p}|\geq 4$, the only proper normal subgroup of $SL_{2}(\mathbb{F}_{p})$ is {±$\{\pm Id\}$ (see for example [30, Theorem 8.3 and Lemma 8.6, p.539]). Thus $\mathrm{SL}_{2}(\mathbb{F}_{p})\subseteq N$. Since the projection over $(\mathbb{F}_{p}^{\times})^{k-1}$ is surjective, the result follows. ∎
@@ Acknowledgements
We are grateful to Andrea Conti, Samuel Le Fourn, Davide Lombardo and Arnaud Plessis for helpful discussions.
We thank the INdAM GNSAGA for its support.
We acknowledge financial support from the CNRS IEA (International Emerging Action) project PAriAIPP (Problèmes sur l’Arithmétique et l’Algèbre des Petits Points).